OpenStudy (anonymous):

Horatio needs to rent a car while on vacation. The cost of the vehicle is $18 per day. He gives the office manager $130 up front. If x represents the number of days and y represents the total cost of the rental, which equation will help horatio determine if he owes the agency money at the end of this trip or if he agency owes him money.

OpenStudy (anonymous):

y=$18x+$130 y=$18x-$130 y=$130-$18 y=$130x+$18

OpenStudy (anonymous):

y=$130x-$18

ganeshie8 (ganeshie8):

okay whats your best guess ?

ganeshie8 (ganeshie8):

Notice that each day rent is 18, so the money he owes grows up 18 per day : 18x

ganeshie8 (ganeshie8):

based on that you can eliminate two options right away

OpenStudy (anonymous):

i think it's y=$18x-$130

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Houston has been hired as a salesman. He is paid a flat rate of $515 each week and earns an additional $17.50 commission for each sale. If his goal is to make $760 in one week, how many sales must Houston make? 43 30 73 14 I picked 14

ganeshie8 (ganeshie8):

looks good ! openstudy is acting up on my end ..

OpenStudy (anonymous):

thanks. mine too

OpenStudy (anonymous):

Joanna pays a flat rate of $74 for her cell phone and is charged $0.08 for every text she sends. Her goal is to spend no more than $96 on her phone bill each month. Which inequality will help her determine the minimum number of texts she sends per month? $74 + $0.08x ? $96 $74 + $0.08x ? $96 $74 + $0.08x < $96 $74 + $0.08x > $96 I picked $74 + $0.08x > $96

OpenStudy (anonymous):

those first one was less than or equal to & the second one was greater than or equal to

ganeshie8 (ganeshie8):

lets say your height is no more than 6 feet

ganeshie8 (ganeshie8):

you represent that using : \[\large \text{height } \le 6\] right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

`no more than 6 feet` means your height could be 6 or less but it can never be more than 6 feet

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

i dont understand why 6 feet is relevant...

ganeshie8 (ganeshie8):

`Her goal is to spend no more than $96`

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so you need to use \(\large \le 96\) inequality here too

OpenStudy (anonymous):

got it!

ganeshie8 (ganeshie8):

you can replace `no more than` with \(\large \le\)

OpenStudy (anonymous):

so it is $74 + 0.08x greater than or equal to 96

ganeshie8 (ganeshie8):

it should be : $74 + 0.08x `less` than or equal to 96

ganeshie8 (ganeshie8):

the left side quantity should be less than or equal to 96

ganeshie8 (ganeshie8):

because her pocket money is only $96

ganeshie8 (ganeshie8):

she cannot pay more than 96

OpenStudy (anonymous):

ooooh okay got it